OpenStudy (anonymous):

Can someone plz tutor me with basic math! i need to review for my test on thursday!

OpenStudy (ash2326):

@vils123 If you need help in a particular question or topic. Post it, We'd be glad to help:)

OpenStudy (anonymous):

I need help with fractions

OpenStudy (ash2326):

Ok, post any question of it

OpenStudy (anonymous):

ok um 5/4 - 5/9

OpenStudy (ash2326):

We have \[\frac{5}{4}-\frac{5}{9}\] Whenever you have to add or subtract two fractions, first notice their denominators If they are same we can just add or subtract, else we'll try to make them same

OpenStudy (ash2326):

Here the denominators are not same 4 and 9 . Can you tell me the LCM of 4 and 9?

OpenStudy (anonymous):

2x2 and 3x3

OpenStudy (ash2326):

These are just factors, tell me the LCM

OpenStudy (anonymous):

i dnt get LCM

OpenStudy (ash2326):

Least Common Muiltiple

OpenStudy (anonymous):

i dont kknow how to do that

OpenStudy (ash2326):

Okay , suppose we have to find LCM of 4 and 6 First step , write factors of 4 and 6 4=2 *2 6=2*3 LCM is the product of all the uncommon factors and the common factor (listed only once) We can notice the common factor is 2 in 4 and 6 , 4 has one more factor 2 and 6 has 3 so LCM is 2*2*3=12 Can you tell the LCM of 4 and 9?

OpenStudy (anonymous):

so how would i get the lcm

OpenStudy (anonymous):

Why don't you go to Khan academy's web site and watch the videos on fractions. http://www.khanacademy.org/math/arithmetic/fractions/v/finding-common-denominators

OpenStudy (anonymous):

i want to work with tutors

OpenStudy (ash2326):

Vils I've shown you how to find LCM, follow the same procedure to find LCM of 4 and 9

OpenStudy (anonymous):

is the lcm 36?

OpenStudy (ash2326):

Yeah correct:D

OpenStudy (anonymous):

so then what

OpenStudy (ash2326):

Now we found the LCM as 36 we have \[\frac{5}{4}-\frac{5}{9}\] Now LCM=>36/4=9 and 36/9=4 So multiply first fraction's numerator and denominator by 9 and second fraction's numerator and denominator by 4 \[\frac {5\times 9} {4 \times 9}-\frac {5 \times 4}{9 \times 4}\] We get \[\frac{45}{36}-\frac{20}{36}\] we now have \[\frac{45-20}{36}\] finally we get \[\frac{25}{36}\]

OpenStudy (anonymous):

what about improper and proper fractions, mixed numbers

OpenStudy (ash2326):

just wait 5 minutes, I'll explain you all

OpenStudy (anonymous):

k

OpenStudy (ash2326):

Hi , sorry kept you waiting:(

OpenStudy (ash2326):

Are you here?

OpenStudy (anonymous):

hi its ok im here

OpenStudy (anonymous):

?

OpenStudy (ash2326):

A proper fraction is a fraction where numerator is less than denominator like 2/3 , 3/4 so on and so forth

OpenStudy (anonymous):

oh how would i know when to reduce?